# Rethinking the Atmospheric Scattering-driven Attention via Channel and Gamma Correction Priors for Low-Light Image Enhancement

### v3

@@ VI Limitation
While our proposed method shows notable improvements in efficiency and performance, a key limitation warrants consideration. As discussed in the methodology section, our approach benefits from the guidance of Channel-Priors and Gamma Correction, which enhances contrast and visual perception. However, gamma correction for brightening with a small value of estimated gamma can lead to defects or distortions in extreme low-light scenarios, even while maintaining realistic exposure. Fig. 10 shows an example of such a limitation. Addressing this issue will require further refinement for brightening and denoising, which we aim to pursue in future work.
@@ VII Conclusion
This work looks deeper at CPGA-Net, utilizing it as an attention mechanism grounded in theoretical formulas. We propose a stacked and modularized attention module to focus on image details. Additionally, we integrate gamma correction into the local branch, creating a Plug-in Attention module for each CP block. This enhancement makes our approach lightweight yet SOTA in performance, maintaining strong efficiency and stable operation across devices with limited computational resources. In the future, we are striving to improve the unsupervised learning of CPGA-Net+ and integrate our approach into HDR imaging and exposure fusion, improving detail preservation in both bright and dark areas by leveraging brightness sensitivity through prior knowledge, such as channel and gamma-correction priors. This will enhance the output’s dynamic range and overall fidelity in real-world applications.
@@ Acknowledgment
This work was supported in part by the Ministry of Science and Technology, Taiwan, under Grant NSTC 113-2221-E-033-055.

### v4

@@ VI Limitation
While our proposed method shows notable improvements in efficiency and performance, a key limitation warrants consideration. As discussed in the methodology section, our approach benefits from the guidance of Channel-Priors and Gamma Correction, which enhances contrast and visual perception. However, gamma correction for brightening with a small value of estimated gamma can lead to defects or distortions in extremely low-light scenarios, even while maintaining realistic exposure. Fig. 11 shows an example of such a limitation. Addressing this issue will require further refinement for brightening and denoising, which we aim to pursue in future work.
@@ VII Conclusion
This work looks deeper at CPGA-Net, utilizing it as an attention mechanism grounded in theoretical formulas. We propose a stacked and modularized attention module to focus on image details and integrate gamma correction into the local branch, creating a Plug-in Attention module for each CP block. Moreover, we reevaluate the systematic design and propose two expanded applications with structural pruning and maximizing potential through theoretical and practical analysis. This enhancement makes our approach lightweight yet SOTA in performance, maintaining strong efficiency and stable operation across devices with limited computational resources. In the future, we are striving to improve the unsupervised learning of CPGA-Net+ and integrate our approach into HDR imaging and exposure fusion, improving detail preservation in both bright and dark areas by leveraging brightness sensitivity through prior knowledge, such as channel and gamma-correction priors. This will enhance the output’s dynamic range and overall fidelity in real-world applications.
@@ Acknowledgment
This work was partly supported by the Ministry of Science and Technology, Taiwan, under Grant NSTC 113-2221-E-033-055.
Figure 12: Visualization of extracted features [8]. (a) Low-light image $L$; (b) A portion of $L$; (c) $\tilde{t}$ (contribution proportion from $L$); (d) An estimated image of $\tilde{A}$; (e) Reconstructed low-light image $R$; (f) Gamma-corrected image $R^{\gamma}$; (g) Enhanced image $\hat{R}$.
Appendix A Revealing the Explainable Mechanisms Behind Degradation
Our method follows a rule-based learning strategy, leveraging the ATSM and the Retinex theory to perform image processing based on physical models. The ATSM simulates the scattering and absorption of light in uneven media, while the Retinex theory, inspired by retinal imaging, models the human visual system. Although these methods provide solid theoretical support from both physical and physiological perspectives, challenges remain in the interpretability of neural networks. While applying the models to deep learning, it is challenging to explain their decision-making mechanisms due to the highly nonlinear nature of deep learning models. Similarly, these aspects are difficult to define. They cannot be easily explained through experience or a basic understanding of the equations derived from the generated features in the neural network. To improve the interpretability of our approach, we use the traditional image degradation model, which offers a more explicit theoretical framework for the deep learning process. This effectively enhances the transparency of the entire neural network, opening the “black box” of image enhancement.
The fundamental image degradation model in matrix form [1] can be defined as:
$\textbf{G}=\textbf{HF}+\textbf{N}$ (18)
Here, the estimated and enhanced image $\hat{R}$ corresponds to $\hat{\textbf{F}}$. Based on the visual observation results from [8], the local branch can be seen as a reconstruction and denoising process for the noise-free or ground-truth reconstructed image $R^{\text{GT}}$, which is equivalent to $\textbf{R}=\textbf{G}-\textbf{N}$, and the global branch can be viewed as an information extraction process for $\textbf{H}^{-1}\textbf{R}$ to lighten the image for a more natural appearance. The characteristic of $\textbf{H}^{-1}$ is extracted by the IAAF module represented by Eq. (5) in the main text. Some relevant images under the current discussion are shown in Fig. 12.
We attempt to further verify our hypotheses for the local branch through least-squares optimization [40]. First, we make the following simple assumptions: we assume that $\hat{A}$ and $L$ can be represented as the noise-free reconstructed image $R^{\text{GT}}$ with additional composite noises that consist of three-channel mixing factors, lighting changes, and inherent thermal noises, which are particularly noticeable in low-light environments, as depicted below:
$\tilde{A}=R^{\text{GT}}+N_{A}$ (20)
$L=R^{\text{GT}}+N_{L}$ (21)
where $L$ is the low-light image, $N_{A}$ and $N_{L}$ denote the noise components associated with $\tilde{A}$ and $L$, respectively. While estimating $N_{A}$, we assume that $N_{L}$ remains fixed once the low-light image $L$ is given. Then, we can define the following residual error $r(\tilde{t},N_{A})$ by combining Eq. (3) in the main text, Eq. (20), and Eq. (21):
$r(\tilde{t},N_{A})=R-R^{\text{GT}}=(1-\tilde{t})N_{A}+\tilde{t}N_{L}$ (22)
We further define a cost function $C(\tilde{t},N_{A})$ for the least-squares optimization process:
$C(\tilde{t},N_{A})=\frac{1}{2}\left[r(\tilde{t},N_{A})\right]^{2}$ (23)
$\frac{\partial C}{\partial N_{A}}=r(\tilde{t},N_{A})\cdot\left(1-\tilde{t}\right)=0$ (25)
Case I: $r(\tilde{t},N_{A})\neq 0$: We have
$(N_{A}-N_{L})=0\quad\text{and}\quad\tilde{t}=1$ (26)
which leads to the solution $N_{A}=N_{L}$, resulting in:
$r_{\text{min}}(\tilde{t},N_{A})=N_{L}$ (27)
However, it results in $R=L$ from Eq. (3) in main text, meaning that no restoration effect is involved at all (identity mapping from the input $L$ to the output $R$).
Case II: $r(\tilde{t},N_{A})=0$: We have
$R-R^{\text{GT}}=0\implies R=R^{\text{GT}}$ (28)
and
Therefore, the optimal solution exists when Eq. (29) holds.